# Diverse Perspectives on AI: Examining People’s Acceptability and Reasoning of Possible AI Use Cases

### v1

@@ Abstract.
In recent years, there has been a growing recognition of the need to incorporate lay-people’s input into the governance and acceptability assessment of AI usage. However, how and why people judge different AI use cases to be acceptable or unacceptable remains under-explored. In this work, we investigate the attitudes and reasons that influence people’s judgments about AI’s development via a survey administered to demographically diverse participants (N=197). We focus on ten distinct professional (e.g., Lawyer AI) and personal (e.g., Digital Medical Advice AI) AI use cases to understand how characteristics of the use cases and the participants’ demographics affect acceptability. We explore the relationships between participants’ judgments and their rationales such as reasoning approaches (cost-benefit reasoning vs. rule-based). Our empirical findings reveal number of factors that influence acceptance such as general negative acceptance and higher disagreement of professional usage over personal, significant influence of demographics factors such as gender, employment, and education as well as AI literacy level, and reasoning patterns such as rule-based reasoning being used more when use case is unacceptable. Based on these findings, we discuss the key implications for soliciting acceptability and reasoning of AI use cases to collaboratively build consensus. Finally, we shed light on how future FAccT researchers and practitioners can better incorporate diverse perspectives from lay people to better develop AI that aligns with public expectations and needs.
1. Introduction
@@
There are growing calls to regulate AI’s development and integration into society (Pistilli et al., 2023). These efforts, as reflected in the EU AI Act (Parliament, 2023), NIST AI Risk Management framework (of Standards and Technology, 2023), and recent U.S. Executive Order (exe, 2023), have resulted in discussions about whether certain AI use cases should be pursued at all. Despite much progress in this area, it is still not clear how to determine which use cases should be pursued or more heavily regulated. Further, little is known about how lay community members, especially those from marginalized groups, feel about the development of specific AI use cases (ada, 2023; Suresh et al., 2024).
One significant challenge when evaluating the acceptability and impact of specific AI use cases111By use cases, we mean specific scenarios, applications, or problems that an AI system is designed to solve or assist within real-world contexts. is that there can be both positive and negative effects, depending on the context(Mun et al., 2024). For instance, while educational AI can provide affordable and accessible personal tutor, it can also lead to over-reliance of students and diminish the goal of education (Times, 2024; Zhai et al., 2024). To develop a generalizable approach to making decisions about AI use cases, ideally we would understand how people resolve these conflicts. More specifically, first, it is essential to understand differences in judgments about acceptability and likely usage across use cases, and how such judgments relate to scenario characteristics. Second, we need better understanding of the personal factors influencing these judgments, especially as they relate to demographic differences (Kingsley et al., 2024). Third, we need to better understand the reasoning strategies participants use when making judgments about AI use cases, and how those strategies do or do not relate to the judgments that are ultimately made. To form governance and policy decisions that anticipate and address disagreements about the development or regulation about specific AI use cases, these understandings are crucial, especially across groups of people with diverse backgrounds, experiences, and familiarity with AI.
To address these needs, in this work we examine how and why lay people judge various AI use cases as acceptable or unacceptable, asking the following research questions:
RQ1
How do judgments of acceptability vary across a set of distinct AI use cases and their characteristics?
RQ2
What attributes or characteristics of people explain the variation in acceptability judgments?
RQ3
How do people reason through acceptability judgments of AI use cases?
@@
To answer these questions, we develop a survey to collect judgments and reasoning processes of 197 demographically diverse participants with varying levels of experience with AI. We ask participants to report whether a certain AI use case should be developed or not, whether they would use such a system, and ask them to provide rationales for their judgment and conditions that would cause them to change their judgments (Figure 1). We examine ten different AI use cases222We focus on text-based, non-embodied, digital systems, and while we do not specifically discuss the AI user and subject, in our use case description, we follow three of the five concepts used in EU AI Act to describe high risk use cases (Golpayegani et al., 2023): the domain, purpose, and capabilities.. To account for differences between sectors or domains, we select use cases in two categories, professional and personal use, and vary them by required entry-level education and EU AI risk level (Table 1).
@@
To meaningfully differentiate and analyze participants’ reasons and reasoning strategies, we borrow concepts from moral psychology and philosophy. We investigate participants’ rationales through two distinct but sometimes overlapping reasoning patterns: cost-benefit reasoning, which assesses expected outcomes (e.g., ”using AI for this task would save time”), and rule-based reasoning, which evaluates the intrinsic values of the action itself (e.g., ”having humans perform this task would be inherently wrong”) (Cushman, 2013; Cheung et al., 2024). We further explore the moral foundations reflected in participants’ reasoning, with moral foundations theory333We used the five foundational dimensions: Care, Fairness, Loyalty, Authority, and Purity. Although these dimensions have been updated to encompass a broader range of values beyond WEIRD (White, Educated, Industrialized, Rich, and Democratic) populations (Atari et al., 2023), we selected this version for survey brevity. (Graham et al., 2011, 2008). Additionally, to understand aspects of AI that raise concerns, we employ three dimensions based on prior studies (Solaiman et al., 2023; Mun et al., 2024): functionality (system capabilities like performance, bias, and privacy), usage (context of system integration, such as supervision, misuse, or unintended use), and societal impact (effects on individuals, communities, and society, such as job loss and over-reliance).
@@
Our empirical results show general higher acceptance of personal use cases over professional. While both categories of use cases show decreased acceptance with increased entry level education and risk, professional use cases display more variability and disagreements across judgments (RQ1). Acceptability significantly varied among demographic groups and levels of AI literacy, with lower acceptability observed particularly among non-male participants and those familiar with AI ethics (RQ2). Finally, our results show varying distribution of reasoning types across acceptability decisions with rule-based reasoning being associated with negative acceptance as well as concern for societal impact. Further qualitative analysis reveal rules such as the need for humanness in certain use cases whether it be for empathy or interaction (RQ3).
@@
Our findings shed novel light onto the diversity of people’s acceptability and reasoning of AI uses in distinct domains and risk levels. We conclude with a discussion highlighting three key implications: first, diverse methodologies are needed to effectively analyze use cases and their characteristics; second, involving diverse stakeholders is crucial for assessing the acceptability of AI applications, particularly in workplaces; and third, further investigation into human reasoning processes about AI, notably rule-based reasoning, is needed to inform consensus-building in policy making.
2. Related Works
In this section, we briefly summarize the background and related work towards assessing acceptability and impact of AI use cases. In each subsection, we highlight how our work extends prior work.
2.1. Assessing Impact of AI
Recent years have witnessed increasing calls from academics (Kieslich et al., 2023; Hecht et al., 2021; Bernstein et al., 2021; Ashurst et al., 2020; Lin et al., 2020; CVPR, Ethics Guidelines, 2023; ACL, Ethic Policy, 2023; ICML, Publication Ethics, 2023; Olteanu et al., 2023; Center for Advanced Study in the Behavioral Sciences, 2020), government (The Ada Lovelace Insitute, 2022; The White House, 2023; National Artificial Intelligence Research Resource Task Force, 2023), civil society (on AI, 2021; Ada Lovelace Institute, 2022; The Ada Lovelace Insitute, 2022; Society, 2023; Metcalf et al., 2021; Reisman et al., 2018), and industry (Microsoft, 2022a, b; Walker and Croak, 2021; OpenAI, 2022; Hecht et al., 2021; Deng et al., 2024) to assess the impact of AI systems designed and developed by AI researchers and practitioners. This effort has particularly highlighted the need to understand the positive impact while grappling with the potential negative impact of integrating AI into certain products and services that affect people’s daily lives (Ashurst et al., 2020; Lin et al., 2020; Hecht et al., 2021). In response, researchers in FAccT, HCI, and AI have developed tools and processes to support AI researchers and practitioners in anticipating the impact of AI systems they developed(Wang et al., 2024; Kieslich et al., 2023; Buçinca et al., 2023; Deng et al., 2024; Weidinger et al., 2022). For example, many have developed AI impact taxonomies or checklists to help developers categorize AI impact(Weidinger et al., 2022; Microsoft, 2022b). Wang et al. and Deng et al. developed tools and templates to support industry AI developers and researchers in assessing the potential negative societal impact of their work, such as job displacement or stereotyping social groups.
However, this prior work primarily focuses on supporting AI experts rather than diverse lay people’s impact assessments of potential AI use cases. Our work extends these prior efforts by understanding diverse (and sometimes conflicting) perspectives on both positive and negative impact of AI use cases from lay people, as a crucial step to complement the AI impact assessments conducted by AI researchers and practitioners.
2.2. Understanding People’s Perceptions of AI Use Cases
Responding to the calls on meaningfully engaging lay people in assessing the impact of specific AI use cases, prior work have started to understand lay people’s perceptions of AI use case (Buçinca et al., 2023; Kieslich et al., 2023; Mun et al., 2024; Kingsley et al., 2024). Among other findings, this prior work revealed a substantial amount of disagreement regarding the desired behavior of AI, primarily due to the subjectivity inherent in certain tasks (e.g., toxicity detection (Sap et al., 2019; Blodgett et al., 2020), image captioning (Zhao et al., 2021)) and ambiguous ethical implication of decisions made by AI for certain tasks (e.g., self driving cars (Awad et al., 2018), medical AI (Chen et al., 2023), predictive analysis (Barocas and Selbst, 2016)). Work done by Mun et al. highlighted that lay people can envision diverse set of harms specific to different AI use cases, complementary to those defined by experts. Another line of work also begins to examine how factors such as demographic backgrounds and previous exposure to discrimination can affect people’s sensitivity towards potential AI harms (Kingsley et al., 2024).
Our work extends this prior work by examined the detailed reasoning processes of lay people regarding the acceptability and the trade-offs between positive and negative impacts of AI use cases. In particular, we draw on model decision theory framework, such as the moral foundations developed by Graham et al., to design a survey flow (See Figure 1 to solicit lay people’s decision-making processes (and potential moral conflicts) when assessing both the benefits and harms of concrete AI use cases.
2.3. Background: Moral Decision Making
Morality, characterized by diverse values across cultures and social groups, aims to suppress selfishness to facilitate social life (Kesebir and Haidt, 2010). To understand decision-making in AI use cases, we draw on moral psychology and dual system theory. We examine two decision-making systems: cost-benefit reasoning, which assesses outcomes and consequences, and rule-based reasoning, focusing on norms, rules, and virtues (Cushman, 2013; Cheung et al., 2024). These correspond to utilitarian reasoning (maximizing good) and deontological reasoning (duties and rights), respectively. Additionally, we apply moral foundations theory (Graham et al., 2008) to identify values and potential moral conflicts in AI development.
3. Study Design and Data Collection
@@
To understand the decision making factors and reasoning patterns of a diverse population regarding AI use cases, we conducted a survey-based study with demographically diverse participants. In this section, we discuss the selection of use cases (§ 3.1), survey design (§ 3.2), and data collection details and participant demographics (§ 3.3).
3.1. Use Cases
@@
To understand how different characteristics of AI use cases can impact judgments and decision making processes, we crafted ten different AI use cases. We first chose two broad application categories frequently mentioned by the public in previous works (Kieslich et al., 2024; Mun et al., 2024): AI in personal, everyday usage where participants could uniformly consider themselves as AI users and AI in professional usage where AI takes on a role thus far done by a human as a profession, e.g., Lawyer AI. We then developed five use cases in the AI in professional usage category that varied according to the level of education required for entry and five use cases in AI in personal category that varied according to the risk level assigned by the EU AI act. All the personal use cases were in the health domain, reflecting a key domain of interest conveyed by lay users in previous works (Mun et al., 2024; Kieslich et al., 2024).
@@
| Professional Use Cases | ||
| Lawyer | Doctoral/Professional Degree | Advises clients on digital legal proceedings/transactions. |
| Elementary School Teacher | Bachelor’s degree | Teaches academic skills at the elementary school level. |
| IT Support Specialist | Some college, no degree | Maintains computer networks and provides technical help. |
| Government Eligibility Interviewer | High school diploma | Determine eligibility for government programs/resources. |
| Telemarketer | No formal education | Solicits donations or orders over the telephone. |
| Personal Use Cases | ||
| Digital Medical Advice | High Risk | Provide medical assessments prior to medical consultations. |
| Customized Lifestyle Coach | High / Limited Risk | Personalized advice for healthy living and wellness. |
| Personal Health Research | Limited Risk | Summarizes research related to personal health issues. |
| Nutrition Optimizer | Limited / Low Risk | Personalize meals and optimize nutritional intake. |
| Flavorful Swaps | Low Risk | Suggest delicious and healthy alternatives food options. |
Professional Use Case Scenarios
@@
For the first area of focus, AI in labor replacement, we collected jobs listed in the U.S. census bureau444https://www.bls.gov/ooh/occupation-finder.htm and sorted them according to entry level education required as stated in the census. We chose education level as it has been tightly linked to socioeconomic and occupational status (Svensson, 2006; Evetts, 2006). We selected jobs that have a large portion of digital or intellectual components with minimal requirement for embodiment resulting in following five professional roles: Lawyer, Elementary school teacher, IT support specialist, Government support eligibility interviewer, and Telemarketer. See Table 1 for further details.
Personal Use Case Scenarios
@@
To understand the acceptability of different health applications in personal and private life, we adapted the descriptions of personal use cases written by participants from prior works (Mun et al., 2024; Kieslich et al., 2024). The research team iteratively refined the descriptions to reflect risk levels according to EU AI Act, and confirmed agreement with categories assigned by GPT-4, following Herdel et al.. See Table 1 for further details.
3.2. Survey Design
@@
Our survey presents participants with five use case descriptions, all from either the professional or personal category (randomly assigned and presented in random order) (see § 3.1 for details). After each description, participants answer: ”Do you think a technology like this should be developed?” (Q1) and then, ”How confident are you in your above answer?” (Q2). They provide open-text rationales by completing the prompt, ”[Use Case] should be developed because…” (Q3), tailored to their Q1 response (”should” for ”Yes” and ”should not” for ”No”). Participants also specify a condition for switching their decision with, ”[Use Case] should not be developed if…” (Q4), adjusted similarly to Q1. Subsequently, they answer, ”If [Use Case] existed, would you use its service?” (Q5) and express confidence with, ”How confident are you in your above answer?” (Q6). Refer to Table 6 in the Appendix for the exact wording of the questions.
@@
Following the main survey, we asked participants questions about their AI literacy and demographics to explore various factors affecting perception of AI acceptance. We adopted a shortened version of AI literacy questionnaires from previous works (Wang et al., 2023; Mun et al., 2024) with four AI literacy aspects, AI awareness, usage, evaluation, and ethics, and two additional questions for generative AI, usage frequency and familiarity with limitations. We collected demographic information of the participants such as race, gender, age, sexual orientation, religion, employment status, income, and level of education. Additionally, we collected information about chronicity, i.e., prolonged experiences of everyday discrimination, of their discrimination experiences (if any) following Kingsley et al..
3.3. Data Collection and Participant Demographics
We used Prolific555https://www.prolific.com to recruit participants. To represent diverse sample, we stratified our recruitment by the ethnicity categories (White, Mixed, Asian, Black and Other) and age (18-48, 49-100) provided by Prolific. We also added criteria for quality such as survey approval rating and number of previous surveys completed. Our study was approved by IRB at our institutions, and we paid 12 USD/hour. Our final sample consisted of 197 participants across two categories, with professional usage assigned to 100 participants and personal to 97. See Appendix A.2 for further details on participants.
4. Analysis Methods

### v2

# Why (not) use AI? Analyzing People’s Reasoning and Conditions for AI Acceptability

@@ Abstract
In recent years, there has been a growing recognition of the need to incorporate lay-people’s input into the governance and acceptability assessment of AI usage. However, how and why people judge acceptability of different AI use cases remains under-explored, despite it being crucial towards understanding and addressing potential sources of disagreement. In this work, we investigate the demographic and reasoning factors that influence people’s judgments about AI’s development via a survey administered to demographically diverse participants (N=197). As a way to probe into these decision factors as well as inherent variations of perceptions across use cases, we consider ten distinct labor-replacement (e.g., Lawyer AI) and personal health (e.g., Digital Medical Advice AI) AI use cases. We explore the relationships between participants’ judgments and their rationales such as reasoning approaches (cost-benefit reasoning vs. rule-based). Our empirical findings reveal a number of factors that influence acceptance. We find lower acceptance of labor-replacement usage over personal health, significant influence of demographics factors such as gender, employment, education, and AI literacy level, and prevalence of rule-based reasoning for unacceptable use cases. Moreover, we observe unified reasoning type (e.g., cost-benefit reasoning) leading to higher agreement. Based on these findings, we discuss the key implications towards understanding and mitigating disagreements on the acceptability of AI use cases to collaboratively build consensus.111We will release data upon acceptance.
1 Introduction
@@
There is a growing call from the public and experts alike to regulate the development and integration of AI into society (Pistilli et al. 2023; McClain 2025). These efforts, as reflected in the EU AI Act (Parliament 2023), NIST AI Risk Management framework (of Standards and Technology 2023), and recent U.S. Executive Order (exe 2023), have resulted in discussions about whether certain AI use cases should be pursued at all. As these high-stakes decisions shape the future of AI, it is essential to equitably determine which AI use cases warrant pursuit despite potential harms, requiring diverse public and expert perspectives. Furthermore, to mitigate potential disagreements, we must understand how individuals make such decisions.
One significant challenge when evaluating the acceptability and impact of AI use cases is that their effects can be simultaneously positive and negative, depending on the context of use, functionality, and broader societal implications (Mun et al. 2024). For instance, while educational AI can provide affordable and accessible personal tutor, it can, at the same time, lead to over-reliance of students and diminish the goal of education (Times 2024; Zhai, Wibowo, and Li 2024). Thus, as AI is applied across increasingly diverse domains, understanding how people make decisions about its use—especially when benefits and harms conflict—becomes critical for anticipating and addressing disagreements about specific use cases.
To tackle this challenge, we investigate the factors and reasoning that shape judgments of AI acceptability. First, we assess how judgments about the acceptability of development and use vary across different AI use cases,222By use cases, we mean specific real-world scenarios or problems that an AI system is designed to address. and how they relate to scenario characteristics (RQ1). Second, we explore personal factors influencing these judgments, especially as they relate to demographic differences (Kingsley et al. 2024) (RQ2). Third, we analyze reasoning strategies participants use when making judgments about AI use cases, and how those strategies do or do not relate to the judgments that are ultimately made (RQ3).
@@
To answer these questions, we develop a survey to collect judgments and reasoning processes of 197 demographically diverse participants with varying levels of experience with AI. We ask participants to report whether a certain AI use case should be developed or not, whether they would use such a system, and ask them to provide rationales for their judgment and conditions that would cause them to change their judgments (Figure 1). We perform a focused investigation of acceptability using ten AI use cases333We focus on text-based, non-embodied, digital systems, and while we do not specifically discuss the AI user and subject, in our use case description, we follow three of the five concepts used in EU AI Act to describe high risk use cases (Golpayegani, Pandit, and Lewis 2023): the domain, purpose, and capabilities. that we systematically select for different risk levels, spanning two highly-discussed domains with ongoing efforts to develop such use cases: personal health and labor replacement (McClain 2025; Kelly 2025; Kolata 2024; Pierson et al. 2025; Rajpurkar et al. 2022; Lee 2024). To understand characteristics of AI use cases that might affect perceptions beyond category, we vary them by required entry-level education and EU AI risk level (Table 1).
@@
We perform a multi-pronged analyses of people’s rationales. Drawing from moral philosophy, we examine participants’ answers using two reasoning patterns: cost-benefit reasoning, which assesses expected outcomes (e.g., “using AI for this task would save time”; akin to utilitarian reasoning), and rule-based reasoning, which evaluates the intrinsic values of the action itself (e.g., “having humans/AI perform this task would be inherently wrong”; akin to deontological reasoning) (Cushman 2013; Cheung, Maier, and Lieder 2024). We then analyze the moral frameworks participants apply, drawing on moral foundations theory (Graham et al. 2011, 2008), to identify the dimensions they prioritize in decision making. Finally, to understand conditions under which participants might flip there decisions, we employ three dimensions based on prior studies (Solaiman et al. 2023; Mun et al. 2024): functionality (system capabilities like performance, bias, and privacy), usage (context of system integration, such as supervision, misuse, or unintended use), and societal impact (effects on individuals, communities, and society, such as job loss and over-reliance).
@@
Our empirical results show general higher acceptance of personal health use cases over labor-replacement. While participants’ acceptability judgments decreased with increased entry-level education and risk for each category respectively, professional use cases display more variability and disagreements across judgments (RQ1). Acceptability significantly varied among demographic groups and levels of AI literacy, with lower acceptability observed particularly among non-male participants and those familiar with AI ethics (RQ2). Finally, our results show varying distribution of reasoning types across acceptability decisions, with rule-based reasoning being associated with negative acceptance and unified reasoning types showing higher agreement. Further qualitative analysis revealed participants’ normative assumptions about AI, humanness, and society—for example, viewing empathy as essential to humanness but lacking in AI (RQ3).
@@
Our findings shed novel light onto the diversity of people’s acceptability and reasoning of AI uses in distinct domains and risk levels. We conclude with a discussion highlighting three key implications for future researchers, practitioners, and policymakers working on advancing ethical and responsible AI development: first, diverse methodologies are needed to effectively analyze use cases and their characteristics; second, involving diverse stakeholders is crucial for assessing the acceptability of AI applications, particularly in workplaces; and third, further investigation into human reasoning processes about AI, notably rule-based reasoning, is needed to inform consensus-building in policy making.
2 Related Works
While there were many efforts towards ethical AI development and deployment by academics (Kieslich, Diakopoulos, and Helberger 2023; Bernstein et al. 2021; Lin et al. 2020), industry (OpenAI 2022; Deng, Barocas, and Vaughan 2024), and government (The White House 2023), they have largely lacked diverse public inputs. To address this gap, many works from both academia and civil society have sought to meaningfully engaging lay people in assessing the impact of specific AI use cases. Prior works have focused on anticipating harms (Buçinca et al. 2023) and impacts (Kieslich, Diakopoulos, and Helberger 2023; ada 2023) through participatory foresight, uncovering diverse, sometimes diverging, viewpoints about AI biases and values (Kingsley et al. 2024; Jakesch et al. 2022; Kapania et al. 2022), and governance efforts of AI (Zhang and Dafoe 2020). However, to the best of our knowledge, only Mun et al. considered development decisions by diverse lay-users with an option of not developing a use cases. Among other findings, these prior works from AIES and broader Responsible AI venues revealed a substantial amount of variations in perceptions primarily among demographic lines (e.g., gender, race, political leaning) regarding the desired behavior of AI.
However, little attention has been given to identifying reasoning of participants over AI use cases. While some works have identified decision variations under ambiguous ethical implications of decisions made by AI for certain tasks (e.g., self driving cars (Awad et al. 2018), medical AI (Chen et al. 2023), predictive analysis (Barocas and Selbst 2016)) and inherent value conflicts (Jakesch et al. 2022), these works have not focused on self-reported reasoning. Thus, our work addresses gap by closely examining the detailed, self-reported reasoning processes of lay people regarding the acceptability of AI use cases without explicitly guiding towards outcome-based (i.e., utilitarian) or value-based (i.e., deontological) reasoning, allowing participants to freely choose and express their deliberation process.
3 Study Design and Data Collection
@@
To answer our research questions on how and why people judge AI use cases as acceptable, we conducted a survey-based study with demographically diverse participants. In this section, we discuss the selection of use cases (§ 3.1), survey design (§ 3.2), and data collection details and participant demographics (§ 3.3).
3.1 Use Cases
@@
To answer RQ1 which examines the impact of different characteristics AI use cases on judgments and decision making processes, we carefully crafted ten different AI use cases as vignettes. We first chose two broad application categories frequently mentioned by the public in previous works (Kieslich, Helberger, and Diakopoulos 2024; Mun et al. 2024): AI in labor-replacement where AI takes on a role in society thus far done by a human as a profession (e.g., Lawyer AI), and AI in personal health, where participants could uniformly consider themselves as AI users. We systematically developed five use cases for each category, varying by required education level for labor-replacement applications and by EU AI Act-assigned risk level for personal health applications.
@@
| Labor-replacement Use | ||
| Lawyer | Doctoral/Prof | Digital legal advice |
| Elementary Teacher | Bachelor | Teaches elementary students |
| IT Support | Some college | Maintains networks; tech support |
| Eligibility Interviewer | High school | Determines benefit eligibility |
| Telemarketer | None | Calls to sell/solicit |
| Peronal Health Use | ||
| Digital Medical Advice | High Risk | Medical advice pre-consultation |
| Lifestyle Coach | High/Limited | Personalized wellness advice |
| Health Research | Limited | Summarizes personal health info |
| Nutrition Optimizer | Lim./Low | Personalized meal/nutrition tips |
| Flavorful Swaps | Low | Suggests healthy food alternatives |
Labor-replacement Use Case Scenarios
@@
For the first area of focus, AI in labor replacement, we collected jobs listed in the U.S. census bureau444https://www.bls.gov/ooh/occupation-finder.htm and sorted them according to entry level education required as stated in the census. We chose education level as it has been tightly linked to socioeconomic and occupational status (Svensson 2006; Evetts 2006). We selected jobs that have a large portion of digital or intellectual components with minimal requirement for embodiment resulting in following five professional roles: Lawyer, Elementary school teacher, IT support specialist, Government support eligibility interviewer, and Telemarketer. See Table 1 for further details.
Personal Health Use Case Scenarios
@@
To understand the acceptability of different health applications in personal and private life, we drew from use cases written by participants from prior works (Mun et al. 2024; Kieslich, Helberger, and Diakopoulos 2024) to systematically craft use cases which varied by risk levels according to EU AI Act. We ensured accurate reflection of the risk levels through iterative refinement of descriptions and agreement with categories assigned by GPT-4, following Herdel et al.. See Table 1 for further details.
3.2 Survey Design
@@
Our survey presents participants with five use case descriptions in random order, all from randomly assigned category, labor-replacement or personal health (see § 3.1 for details). After each description, participants answer: “Do you think a technology like this should be developed?” (Q1) and then, “How confident are you in your above answer?” (Q2). To allow for examining of their reasoning, participants then provide open-text rationales by finishing the sentence, “[Use Case] should [not] be developed because…” (Q3), adjusted dynamically depending on their answer to Q1. Participants also described that they would switch their opinion on acceptability of development of the use case: however, “[Use Case] should [not] be developed if…” (Q4; also dynamically rephrased based on Q1 answer). Subsequently, they answer, “If [Use Case] existed, would you use its service?” (Q5) and express confidence with, “How confident are you in your above answer?” (Q6). Refer to Table 6 in the Appendix for the exact wording of the questions.
@@
Following the main survey, we asked participants questions about their AI literacy level and demographics to explore various factors affecting perception of AI acceptance (RQ2). We adopted a shortened version of AI literacy questionnaires from previous works (Wang, Rau, and Yuan 2023; Mun et al. 2024) with four AI literacy aspects, AI awareness, usage, evaluation, and ethics, and two additional questions for generative AI, usage frequency and familiarity with limitations. We collected demographic information of the participants such as race, gender, age, sexual orientation, religion, employment status, income, and level of education; see Appendix 7 for detailed list of questions. Additionally, we collected information about discrimination chronicity, i.e., prolonged experiences of everyday discrimination, of their discrimination experiences (if any) following Kingsley et al..
3.3 Data Collection and Participant Demographics
We used Prolific555https://www.prolific.com to recruit participants. To represent diverse sample, we stratified our recruitment by the ethnicity categories (White, Mixed, Asian, Black, and Other) and age (18-48, 49-100) as provided by Prolific. We also added criteria for quality such as survey approval rating and number of previous surveys completed. Our study was approved by IRB at our institutions, and we paid 12 USD/hour. Our final sample consisted of 197 participants across two categories, with professional usage assigned to 100 participants and personal to 97. See Appendix A.3 for further details on participants.
4 Acceptability & Reasoning
Analysis Methods